OpenStudy (anonymous):

Look at the pattern shown below. What number is missing from the pattern? 20, 22, 26, 34, __,82 50 Not Here 52 38 42

OpenStudy (anonymous):

Number 50 is mising ;)

OpenStudy (anonymous):

Explanation: difference between numbers is 2,4,8,16,32 20 + 2 = 22 + 4 = 26 + 8 = 34 + 16 = 50 + 32 = 82